# To Rely or Not to Rely? Evaluating Interventions for Appropriate Reliance on Large Language Models

### v1

@@ 5.1. Key Findings
The increase in advice under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b). This suggests that for the LSAT task, static Reliance Disclaimers may have added just enough friction to dissuade habitual over-reliance. However, advice with other interventions, like Implicit Answer in Numerical Estimation, may have been perceived as requiring more effort to use than it was worth. Future work should aim to quantify the comprehension difficulty of the interventions and the tasks to understand how they relate to reliance behaviours. In implementing interventions, designers should consider the targeted use case and develop cognitive forcing techniques that do not increase burden substantially.
Confidence When Using LLMs is Poorly Calibrated. We examine how participants’ confidence in their answers change after seeing the LLM’s advice. Ideally, even with incorrect reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice. However, we largely observe the opposite or inconclusive effects, where wrong reliance results in higher gains in confidence. This raises particularly problematic implications, suggesting that individuals were overall uncalibrated when receiving assistance from LLMs, corroborating prior findings (Zhang, 2023; Si et al., 2023)
@@ 5.3. Limitations
Engagement of Participants in Tasks. The generalizability of our results is primarily limited by our design of a lab experiment with crowdworkers recruited from Prolific. Since participants knew that they were participating in an experiment, they may have inferred that they were being evaluated on LLM reliance and become more cautious about their reliance behaviours. Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to “hedge”; to maximize their chances of earning some bonus payment by choosing different answers at each stage. Our two-stage question answering set-up also emulates a specific type of decision-making where users are already engaged in the task. In fact, having users answer questions independently before advice can be considered a form of cognitive forcing even without additional interventions (Ashktorab et al., 2024). As LLM usage becomes more common, it is expected that many users will bypass this critical engagement and rely on LLMs by default, which can reflect in reliance outcomes differently.
@@
Realism of Interacting with the LLM. Due to the nature of the survey study, our participants were unable to interactively ask follow-up questions or use sanity checks to verify the model’s outputs and prevent consequential errors. To ensure clean reliance metrics, they saw fixed advice and were asked not to consult external resources. Although we used direct outputs generated through tailored LLM prompts whenever possible, we did apply manual adjustments in some interventions and artificially manipulated the recommendation in the LLM’s advice. The distribution of advice quality was held constant regardless of the LLM’s true tendencies when responding to the question to prevent question difficulty from being a confound. We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude555https://claude.ai/, Copilot666https://copilot.microsoft.com/, Gemini777https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour. We also did not explore varying the system prompts to steer the LLM into embodying different personas. Furthermore, the framing of the LLM (e.g., ChatGPT vs. a generic chatbot) can prime user expectations of the quality and accuracy of the responses.
@@ Appendix A Task Question Examples
Figures A.1 and A.2 show screenshots LSAT Logical Reasoning tasks questions. Participants answer independently in Stage 1, and then re-answer the same question with randomly sampled LLM advice in Stage 2.
Figures A.3 and A.4 show screenshots of Numerical Estimation task questions, with the same two-stage design.
@@ Appendix D Confidence Analysis for Other WoA Thresholds
The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.8 and D.9.
@@ Appendix E Effect of Prior LLM Experience
As an additional analysis, we examine whether participants’ baseline frequencies of using LLM tools affected their behaviours and perceptions. The participants reported their prior experience with LLMs in the survey as never tried, almost never but have tried, monthly, weekly, and daily, which we map into ordinal values from 0 to 4. When controlling for condition, a partial Spearman correlation reveals no significant correlations between LLM usage and reliance on the advice for either task, LSAT ( for a binary reliance variable) or Numerical Estimation ( for the continuous WoA variable). Nor are there correlations with the changes in confidence for LSAT () and Numerical Estimation (). However, there is significant positive association between frequency of usage and all three categories of post-task perception: confidence (), accuracy (), and helpfulness (). This suggests that while people who were more familiar and open to using LLM tools generally hold higher view of ChatGPT’s advice (across all conditions), it did not impact actual rates of reliance in the task.

### v2

@@ 5.1. Key Findings
The increase in under-reliance supports the cost-benefit framework from explainable AI literature, which shows that users will not engage with explanations if they are more cognitively complex than the task itself (Vasconcelos et al., 2023b). This suggests that for the LSAT task, static Reliance Disclaimers may have added just enough friction to dissuade habitual over-reliance. However, advice with other interventions, like Implicit Answer in Numerical Estimation, may have been perceived as requiring more effort to use than it was worth. Future work should aim to quantify the comprehension difficulty of the interventions and the tasks to understand how they relate to reliance behaviours. In implementing interventions, designers should consider the targeted use case and develop cognitive forcing techniques that do not increase burden substantially.
Confidence When Using LLMs is Poorly Calibrated. We examine how participants’ confidence in their answers change after seeing the LLM’s advice. Ideally, even with wrong reliance behaviour, confidence should be lower when individuals are internally uncertain about their decision, such as when they are skeptical of the LLM’s advice. However, we largely observe the opposite or inconclusive effects, where wrong reliance results in higher gains in confidence. This raises particularly problematic implications, suggesting that individuals were overall uncalibrated when receiving assistance from LLMs, corroborating prior findings (Zhang, 2023; Si et al., 2023)
@@ 5.3. Limitations
Engagement of Participants in Tasks. The generalizability of our results is primarily limited by our design of a lab experiment with crowdworkers recruited from Prolific. Since participants knew that they were participating in an experiment, they may have inferred that they were being evaluated on LLM reliance and become more cautious about their reliance behaviours. Although we tried to emulate the risks and benefits of real-world decision-making by rewarding participants for correctness, this may have encouraged participants to ‘hedge’; to maximize their chances of earning some bonus payment by choosing different answers at each stage. Our two-stage question answering set-up also emulates a specific type of decision-making where users are already engaged in the task. In fact, having users answer questions independently before advice can be considered a form of cognitive forcing even without additional interventions (Ashktorab et al., 2024). As LLM usage becomes more common, it is expected that many users will bypass this critical engagement and rely on LLMs by default, which can reflect in reliance outcomes differently.
@@
Realism of Interacting with the LLM. Due to the nature of the survey study, our participants were unable to interactively ask follow-up questions or use sanity checks to verify the model’s outputs and prevent consequential errors. To ensure clean reliance metrics, they saw fixed advice and were asked not to consult external resources. Although we used direct outputs generated through tailored LLM prompts whenever possible, we did apply manual adjustments in some interventions and artificially manipulated the recommendation in the LLM’s advice. The distribution of advice quality was held constant regardless of the LLM’s true tendencies when responding to the question to prevent question difficulty from being a confound. We only explore one specific LLM, GPT-4o from OpenAI, so chatbots like Claude777https://claude.ai/, Copilot888https://copilot.microsoft.com/, Gemini999https://gemini.google.com/, and other open-source models may reflect different conversational styles that can implicitly affect user behaviour. We also did not explore varying the system prompts to steer the LLM into embodying different personas. Furthermore, the framing of the LLM (e.g., ChatGPT vs. a generic chatbot) can prime user expectations of the quality and accuracy of the responses.
@@ Appendix A Task Question Examples
Figure 1(b) shows an example of the LSAT Logical Reasoning task question. Participants answer independently in Stage 1, and then re-answer the same question with randomly sampled LLM advice in Stage 2.
The “Road” LSAT Logical Reasoning question is depicted via two figures labelled (a) and (b). Screenshot (a) depicts Stage 1. A multiple-choice question’s text presents a paragraph making an argument, then asks “Which of the following, if true, most strengthens the argument in the paragraph above?” Options A-E are shown below. Screenshot (b) depicts Stage 2. Text at the top of this screenshot says “Now you have access to ChatGPT’s advice. Will you change your answer? Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same question from above is repeated. Below we see sample advice from ChatGPT followed by “Please enter what you believe is the correct answer”. Below that, the multiple choice options are displayed again.
Figures A.2 and A.3 show screenshots of Numerical Estimation task questions, with the same two-stage design.
The Chickpea Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b). Screenshot (a) depicts Stage 1. Question text asks “How many chickpeas are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of chickpeas. Below we see “Please enter a number representing the number of chickpeas:” followed by a text box. Screenshot (b) depicts Stage 2. Text at the top of this screenshot says “Now you have access to ChatGPT’s advice. Will you change your answer? Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the chickpea jar are shown. Below, we see the participant’s Stage 1 answer. Below this, we see sample advice from ChatGPT. Below that, we see “Please enter a number representing the number of chickpeas:” followed by a text box.
The Jellybean Numerical Estimation question is depicted via two Qualtrics screenshots labelled (a) and (b). Screenshot (a) depicts Stage 1. Question text asks “How many jellybeans are in this glass jar?” Two pictures are displayed: a bottom-up view and a side view of the same jar of jellybeans. Below we see “Please enter a number representing the number of jellybeans:” followed by a text box. Screenshot (b) depicts Stage 2. Text at the top of this screenshot says “Now you have access to ChatGPT’s advice. Will you change your answer? Note that this is ChatGPT’s answer to the question only; it does not know your answer.” The same two pictures of the jellybean jar are shown. Below we see the participant’s Stage 1 answer. Below this, we see sample advice from ChatGPT. Below that, we see “Please enter a number representing the number of jellybeans:” followed by a text box.
@@ Appendix B Control and Intervention Advice Examples
Figures 1(d) and B.2 show sample ChatGPT advice for all four conditions for a LSAT Logical Reasoning question and a Numerical Estimation, respectively.
Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample LSAT question for each condition. Screenshot (a) depicts Control condition advice, which recommends option B with reasoning. Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink. Screenshot (d) depicts Implicit Answer advice, which has reasoning for why Option B is correct–without explicitly referring to Option B.
Four Qualtrics screenshots labelled (a), (b), (c), and (d) with sample advice for the sample NE question for each condition. Screenshot (a) depicts Control condition advice, which has a 4-step answer. Step 1 estimates chickpea dimensions. Step 2 makes assumptions about jar dimensions and calculates its volume. Step 3 estimates chickpea volume. Step 4 divides jar volume by chickpea volume to get an estimate for total chickpeas, concluding that there are approximately 887 chickpeas. Screenshot (b) depicts Reliance Disclaimer advice, which is identical to Control condition advice but followed with bolded text saying “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.” Screenshot (c) depicts Uncertainty Highlighting advice, which is identical to Control condition advice but with some text that is highlighted pink and dark pink. Screenshot (d) depicts Implicit Answer advice, which is identical to Control condition advice but with the final step explained rather than shown.
@@ Appendix C Levels of Advice in Numerical Estimation
Two figures labelled (a) and (b) depicting the distribution of participant estimations and LLM advice levels for the Numerical Estimation Task, separately for all four conditions. Subfigure (a) is the chickpea estimation question and subfigure (b) is the jelly bean estimation question.
@@ Appendix D Confidence Analysis for Other WoA Thresholds
The confidence calibration graphs for Numerical Estimation at the other WoA thresholds (0.01 and 0.99) are listed here as Figure D.1 and D.2.
NE confidence changes by reliance behaviour, with a WoA threshold of 0.01. Two double bar graphs labelled (a) and (b). Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99. Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition. Average change in confidence is always positive. For each category, there is no statistically significant difference between average change in confidence after wrong vs. correct reliance behaviour. Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2. Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect). Both are grouped by condition. The average change in confidence is always positive. There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
NE confidence changes by reliance behaviour, with a WoA threshold of 0.99. Two double bar graphs labelled (a) and (b). Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.99. Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition. Average change in confidence is always positive. Only Implicit Answer has statistically different changes in confidence for right vs. wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour. Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2. Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect). Both are grouped by condition. The average change in confidence is always positive. There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
@@ Appendix E Effect of Prior LLM Experience
As an additional analysis, we examine whether participants’ baseline frequencies of using LLM tools affected their behaviours and perceptions. The participants reported their prior experience with LLMs in the survey as never tried, almost never but have tried, monthly, weekly, and daily, which we map into ordinal values from 0 to 4. When controlling for condition, a partial Spearman correlation reveals no significant correlations between LLM usage and reliance on the advice for either task, LSAT ( for a binary reliance variable) or Numerical Estimation ( for the continuous WoA variable). Nor are there correlations with the changes in confidence for LSAT () and Numerical Estimation (). However, there is significant positive association between frequency of usage and all three categories of post-task perceptions of the LLM: Confidence (), Accuracy (), and Helpfulness (). This suggests that while people who were more familiar and open to using LLM tools generally hold higher view of ChatGPT’s advice (across all conditions), it did not impact actual rates of reliance in the task.